OpenStudy (anonymous):

Subtract. Simplify if possible 3-y/y-8 - 2y-5/8-y=

OpenStudy (anonymous):

\[\frac{3-y}{y-8} - \frac{2y-5}{8-y}=\]

hero (hero):

\[\frac{3-y}{y-8} + \frac{2y-5}{y-8}\]

hero (hero):

That's what it simplifies to, then you can add them now.

hero (hero):

\[\frac{y-2}{y-8}\]